Events A and B are disjoint.P(A) =40/61 and P(B) = 2/61/
Find P(A or B)

Thank you for your answers. Answers Preferred from Checkley77,Solver 91311,Stanbon,or Fombitz,Thanks.

Answer by stanbon(75887)

Events A and B are disjoint.P(A) =40/61 and P(B) = 2/61/
Find P(A or B)
------------------------
Formula: P(A or B) = P(A) + P(B) - P(A and B)
But A and B are disjoint so P(A and B) = 0
So, P(A or B) = P(A) + P(B) = (40/61) + 2/61 = 42/61
====================
Cheers,
Stan H.